### v1

# An alternative GPU acceleration for a pseudopotential plane-waves density functional theory code with applications to metallic systems

Abstract
We present an alternative GPU acceleration for plane waves pseudopotentials electronic structure codes designed for systems that have small unit cells but require a large number of k points to sample the Brillouin zone as happens, for instance, in metals. We discuss the diagonalization of the Kohn and Sham equations and the solution of the linear system derived in density functional perturbation theory. Both problems take advantage from a rewriting of the routine that applies the Hamiltonian to the Bloch wave-functions to work simultaneously (in parallel on the GPU threads) on the wave-functions with different wave-vectors k, as many as allowed by the GPU memory. Our implementation is written in CUDA Fortran and makes extensive use of kernel routines that run on the GPU (GLOBAL routines) or can be called from inside the GPU threads (DEVICE routines). We compare our method with the CPUs only calculation and with the approach currently implemented in Quantum ESPRESSO that uses GPU accelerated libraries for the FFT and for the linear algebra tasks such as the matrix-matrix multiplications as well as OpenACC directives for loop parallelization. We show in a realistic example that our method can give a significant improvement in the cases for which it has been designed.
keywords:
Electronic structure; Lattice dynamics; Metals; GPU; Density functional perturbation theory.
journal: Computer Physics Communications
1 Introduction
Density functional theory [1] (DFT) and the availability of more and more powerful computers has made the study of material properties from first-principles a well established reality. Several tools have been refined over the years to solve the one electron Kohn and Sham equations that derive from DFT, [2] the most widespread being based on a plane waves basis and pseudopotentials. Well tested, freely available [3, 4, 5] or commercial [6, 7] packages implement the theory and allow the calculation of material properties.
In the last ten years, high performance computers aiming to reach the exaflops ($10^{18}$ floating point operations per second) switched to a hybrid technology in which the graphic processing units (GPUs) support the central processing units (CPUs) in the floating point operations. In theory, the GPUs can deliver one or two orders of magnitude more flops than the CPUs themselves, and to harness this power many electronic structure groups are modifying their codes to run on the GPUs[8, 9, 10, 11, 12, 13]
Extensions of common programming languages such as C/C++ or Fortran have provided commands to allocate data on the GPU, to move data from the CPU to the GPU and vice versa, and to perform calculations on these data with the GPU. CUDA Fortran commands, [14] declarations, and compiler directives and OpenACC compiler directives [15] are two of the most commonly used Fortran extensions. Recently also applications based on openMP started to appear in the literature. [16, 17] Actually OpenACC and OpenMP have the additional benefit of being transferable to GPUs architectures of different vendors such as NVIDIA, AMD, or Intel and sometimes are preferred to CUDA Fortran that is limited to NVIDIA GPUs.
So far, in several plane-waves pseudopotentials and quantum chemistry codes, the GPUs have been exploited by allocating the variables on the GPU and by substituting the calls to linear algebra and fast Fourier transform (FFTs) routines with calls to optimized library routines (such as cuBlas[18] cuSolver[19] cuFFT[20] and MAGMA [21]) developed by the GPUs vendors and capable to run on the GPU[22, 9, 16, 23, 24, 25] Sometimes the routines of these libraries have the same names and arguments of the corresponding CPU libraries and it suffices to allocate the variables on the GPU to call the GPU routines with minimal changes to the underlying codes and algorithms. Single loops using variables allocated on the GPUs can also be accelerated by compiler directives.
In Quantum ESPRESSO [3, 4] work on this kind of acceleration started more than ten years ago [26] and has been improved over the years [27] leading to a well tested package. [9] Accelerations of 2X or higher with respect to the CPU are often found in pseudopotential plane waves codes that adopt this approach. However, test systems are usually big supercells with many atoms for which the time spent to make calculations on the GPU is larger than the time needed to transfer data from the CPU to the GPU, while small size systems are left out from these tests. For some applications, metallic systems with small unit cells need tens or hundreds thousands k points to sample the Fermi surface. [28, 29, 30, 31, 32, 33, 34, 35, 36] The calculation of the phonon dispersions of these systems for many geometries as required for thermodynamic calculations is a problem that could take advantage from the new supercomputers, but for small systems we found that the use of the GPUs with the present codes is not always convenient and sometimes it can also slow down the calculation with respect to the CPUs alone.
We have therefore tried to improve the situation and in this paper we present the solution that we have found: an alternative approach to the acceleration of the pseudopotentials plane waves codes that is useful to deal with metallic system when there are many k points. We load on the GPU many wave-functions (i.e. k points), all the available ones if the GPU memory is large enough or as many as possible until there is free GPU memory. Then we make the calculations simultaneously on all these data (application of the Hamiltonian to the wave-functions) with each GPU thread working on a single wave-function or on a part of it. To obtain the precise control of the GPU threads that is needed we wrote a set of kernel functions (called GLOBAL in the CUDA language) that implement the theory and run on the GPU. These kernel functions need to call linear algebra and FFT library functions from inside the GPU threads. Unfortunately, libraries such as cuFFT, cuSolver, or MAGMA, which are called from the CPU and automatically control the number of threads, are not suited for our algorithm. We need functions that can be called from inside the GPU threads (DEVICE functions in the CUDA language). We only find the C++ library cuFFTDx [37] that implements such functionalities, but it does not provide a FORTRAN interface so far. For the moment, we transformed to the DEVICE form the FORTRAN sources of fftpack5.1 and of selected LAPACK routines. Finally, we obtained a code significantly faster than the standard one for small systems with many k points.
We start with a brief introduction of the main equations that are solved in a plane-waves pseudopotential code. We stress in particular the algorithms that are relevant for the following discussion, neglecting the parts that have not changed or that are still calculated on the CPU. We then discuss how, in our method, the different parts of the code have been accelerated on the GPU. Finally, we present a test of our implementation and compare the times required by our approach with those taken by the CPUs only calculations and by the previously available GPU implementation.
2 Theory
The solutions of the Kohn and Sham (KS) equations minimize the DFT total energy. These equations are an eigenvalue problem for norm conserving pseudopotentials, [38] and a generalized eigenvalue problem for ultrasoft [39] or projector-augmented wave (PAW) pseudopotentials. [40, 41] For periodic solids they can be written as:
${H}_{KS}\psi_{{\bf k}\nu}=\varepsilon_{{\bf k}\nu}S\psi_{{\bf k}\nu},$ (1)
where ${\bf k}$ is a wave vector and $\nu$ is a band index. $H_{KS}$ is the Kohn and Sham Hamiltonian and $S$ is the overlap matrix. We are interested in finding the lowest $N_{b}$ (number of bands) eigenvalues and eigenvectors of these equations. The Kohn and Sham Hamiltonian depends itself from a potential that is calculated from the charge density (that also depends on the wavefunctions). It is possible to solve this problem by a self-consistent procedure in which the wavefunctions are first calculated with an approximate potential. Then these wavefunctions are used to recompute the charge density and a new potential. The latter is mixed with the potential of the previous iterations and the procedure is repeated until one reaches self-consistency. At each step of the procedure however one has to diagonalize a fixed Hamiltonian which is progressively improved.
In the standard algorithm the problem is solved sequentially for each ${\bf k}$ vector and the charge density is computed at the end when all wave-functions are available.
2.1 Davidson algorithm
There are several algorithms currently implemented in electronic structure codes to find the eigenvalues and eigenfunctions in Eq. 1, but here we limit the discussion to the Davidson algorithm. [42] In this algorithm an initial set of $N_{b}$ functions $|\phi_{i}^{(n)}\rangle$ are progressively improved by enlarging the set applying $H_{KS}-\varepsilon_{i}S$ and solving the generalized eigenvalue problem $\tilde{H}_{ij}-\varepsilon\tilde{S}_{ij}$ on the basis formed by the original and the newly calculated vectors. A standard software library for numerical linear algebra, such as LAPACK[43] is employed for the diagonalization. The algoritm is the following and has to be repeated for each k point:
  • 1.
    Given $N_{b}$ trial eigenpairs: $\quad\left\{\left|\phi_{i}^{(n)}\right\rangle,\varepsilon_{i}^{(n)}\right\}$ of the reduced Hamiltonian calculate:
    $\tilde{H}_{ij}=\left\langle\phi_{i}^{(n)}\left|H_{KS}\right|\phi_{j}^{(n)}\right\rangle,\quad\tilde{S}_{ij}=\left\langle\phi_{i}^{(n)}|S|\phi_{j}^{(n)}\right\rangle.$ (2)
  • 2.
    Build the correction vectors $\left|\tilde{\phi}_{i}^{(n)}\right\rangle$:
    $\left|\tilde{\phi}_{i}^{(n)}\right\rangle=\left(H_{diag}-\varepsilon_{i}^{(n)}S_{diag}\right)^{-1}\left(H_{KS}-\varepsilon_{i}^{(n)}S\right)\left|\phi_{i}^{(n)}\right\rangle,$ (3)
    where $H_{diag}$ and $S_{diag}$ are the diagonal elements of $H_{KS}$ and $S$ in the plane waves representation.
  • 3.
    Normalize the correction vectors:
    $|\tilde{\phi}_{i}^{(n)}\rangle={|\tilde{\phi}_{i}^{(n)}\rangle\over\sqrt{\langle\tilde{\phi}_{i}^{(n)}|\tilde{\phi}_{i}^{(n)}\rangle}}.$ (4)
  • 4.
    Build an extended reduced Hamiltonian and overlap matrix:
    $\tilde{H}_{ij}=\left\langle\phi_{i}^{(n)}/\tilde{\phi}_{i}^{(n)}\left|H_{KS}\right|\phi_{j}^{(n)}/\tilde{\phi}_{j}^{(n)}\right\rangle,\quad\tilde{S}_{ij}=\left\langle\phi_{i}^{(n)}/\tilde{\phi}_{i}^{(n)}|S|\phi_{j}^{(n)}/\tilde{\phi}_{j}^{(n)}\right\rangle.$ (5)
  • 1.
    Set $N_{base}$ equal to the number of basis vector. Diagonalize the small $N_{base}\times N_{base}$ reduced Hamiltonian to get the new estimate for the lowest $N_{b}$ eigenpairs:
    $(\tilde{H}-\varepsilon\tilde{S})v=0\quad\\ \longrightarrow\quad\left\{\left|\phi_{i}^{(n+1)}\right\rangle,\varepsilon_{i}^{(n+1)}\right\}.$ (6)
  • 2.
    Calculate $|\tilde{\phi_{i}}^{(n+1)}\rangle$ for all $i$ for which $\left|\varepsilon_{i}^{(n+1)}-\varepsilon_{i}^{(n)}\right|>\varepsilon_{th}$ where $\varepsilon_{th}$ is the accuracy required for the eigenvalues and call $N_{nc}$ the number of new vectors.
  • 3.
    If $N_{nc}>0$ repeat with the basis $|\phi_{i}^{(n)}/\tilde{\phi_{i}}^{(n)}/\tilde{\phi_{i}}^{(n+1)}\rangle$ of size $N_{base}+N_{nc}$ and continue with progressively larger basis. When the size of the basis becomes too large for the allocated memory instead of adding $\quad\left\{\left|\tilde{\phi_{i}}^{(n+1)}\right\rangle\right\}$ to the basis, restart with $\quad\left\{\left|\phi_{i}^{(n+1)}\right\rangle,\varepsilon_{i}^{(n+1)}\right\}$. If $N_{nc}=0$ exit with eigenpairs $\quad\left\{\left|\phi_{i}^{(n+1)}\right\rangle,\varepsilon_{i}^{(n+1)}\right\}$
The most time consuming step of this algorithm is the application of the operators $H_{KS}$ and $S$ to the wave-functions as discussed in the next subsection.
2.2 Application of Hamiltonian
The KS Hamiltonian can be written as: [44]
${H_{KS}}\psi_{{\bf k}\nu}=\underbrace{-\frac{1}{2}\nabla^{2}\psi_{{\bf k}\nu}}_{kinetic\ energy}+\underbrace{V_{eff}\psi_{{\bf k}\nu}}_{local\ energy}+\underbrace{V_{NL}\psi_{{\bf k}\nu}}_{non-local\ energy},$ (7)
where the effective potentials is the sum of the local, Hartree, and exchange and correlation potentials:
$V_{eff}=V_{loc}+V_{H}+V_{XC},$ (8)
while the nonlocal pseudopotential is defined in term of the projector functions $\left|\beta^{I}_{m}\right\rangle$ and pseudopotential coefficients $D^{I}_{mn}$[38, 39]
$V_{NL}\left|\psi_{{\bf k}\nu}\right\rangle=\sum_{Imn}{D^{I}_{mn}}{\left|\beta^{I}_{m}\right\rangle}{\left\langle\beta^{I}_{n}\right|\psi_{{\bf k}\nu}\big{\rangle}}.$ (9)
Here $I$ indicates the different atoms in the solid and $m$ and $n$ run on all the $\beta^{I}_{m}$ functions of a given atom.
The overlap matrix can be calculated in a similar way: [39]
$S\left|\psi_{{\bf k}\nu}\right\rangle=\left|\psi_{{\bf k}\nu}\right\rangle+\sum_{mn}q^{I}_{mn}\left|\beta^{I}_{m}\right\rangle\left\langle\beta^{I}_{n}\big{|}\psi_{{\bf k}\nu}\right\rangle,$ (10)
where the coefficients $q^{I}_{mn}$ are defined together with the pseudopotential.
2.2.1 Kinetic energy
The kinetic energy is calculated in reciprocal space. Using the Bloch theorem we write the Bloch wave-functions as:
$\psi_{{\bf k}\nu}({\bf r})=e^{i{\bf k}{\bf r}}u_{{\bf k}\nu}({\bf r})=\frac{1}{\sqrt{V}}\sum_{{\bf G}}C_{{\bf k}+{\bf G}\nu}e^{i({\bf k}+{\bf G}){\bf r}},$ (11)
where $u_{{\bf k}\nu}({\bf r})$ is a lattice periodic function expanded in plane waves (here $V$ is the volume of the solid) and the sum is over the reciprocal lattice vectors contained into a sphere defined by the relationship:
${1\over 2}|{\bf k}+{\bf G}|^{2}<E_{cut},$ (12)
where $E_{cut}$ is the kinetic energy cut-off. Then we have:
$-{1\over 2}\nabla^{2}\psi_{{\bf k}\nu}({\bf r})={1\over\sqrt{V}}\sum_{{\bf G}}C^{\prime}_{{\bf k}+{\bf G}\nu}e^{i({\bf k}+{\bf G}){\bf r}},$ (13)
where:
$C^{\prime}_{{\bf k}+{\bf G}\nu}={1\over 2}|{\bf k}+{\bf G}|^{2}C_{{\bf k}+{\bf G}\nu}.$ (14)
2.2.2 Local potential
The fast Fourier transform (FFT) transforms functions in real space into reciprocal space and the inverse FFT makes the inverse transformation.
From the coefficients $C_{{\bf k}+{\bf G}\nu}$, applying an inverse FFT we obtain the Bloch function in real space (up to a factor $1/\sqrt{V}$):
$C_{{\bf k}+{\bf G}\nu}\xrightarrow{FFT^{-1}}u_{{\bf k}\nu}({\bf r})=\sum_{{\bf G}}C_{{\bf k}+{\bf G}\nu}e^{i{\bf G}{\bf r}}.$ (15)
The effective potential is applied in real space as:
$u^{\prime}_{{\bf k}\nu}({\bf r})=V_{eff}({\bf r})u_{{\bf k}\nu}({\bf r}),$ (16)
and a final FFT computes the plane wave expansion of $u^{\prime}_{{\bf k}\nu}({\bf r})$:
$u^{\prime}_{{\bf k}\nu}({\bf r})\xrightarrow{{FFT}}C^{\prime}_{{\bf k}+{\bf G}\nu}={1\over N_{\bf r}}\sum_{{\bf r}}u^{\prime}_{{\bf k}\nu}({\bf r})e^{-i{\bf G}{\bf r}},$ (17)
where $N_{\bf r}$ is the number of points of the FFT grid (see below).
The actual calculation of $V_{eff}$ requires the calculation of the charge density in terms of the wave-functions $\psi_{{\bf k}\nu}$. However since we have not modified this part of the calculation we do not discuss it in detail. We assume only to have a function $V_{eff}$ defined in the points of the FFT grid ${\bf r}$.
2.2.3 Non local pseudopotential and overlap matrix
The application of the non local potential needs three matrix-matrix multiplications:
$\lambda^{I}_{n{\bf k}\nu}=\left\langle\beta^{I}_{n}\big{|}\psi_{{\bf k}\nu}\right\rangle=\sum_{\bf G}\beta^{I}_{n}({\bf k}+{\bf G})^{*}C_{{\bf k}+{\bf G}\nu},$ (18)
$\gamma^{I}_{m{\bf k}\nu}=\sum_{n}{D^{I}_{mn}}\lambda^{I}_{n{\bf k}\nu},$ (19)
$V_{NL}\left|\psi_{{\bf k}\nu}\right\rangle=\sum_{Im}\gamma^{I}_{m{\bf k}\nu}\left|\beta^{I}_{m}\right\rangle.$ (20)
Similarly, the application of the $S$ matrix is:
$\delta^{I}_{m{\bf k}\nu}=\sum_{n}{q^{I}_{mn}}\lambda^{I}_{n{\bf k}\nu},$ (21)
$S\left|\psi_{{\bf k}\nu}\right\rangle=\left|\psi_{{\bf k}\nu}\right\rangle+\sum_{Im}\delta^{I}_{m{\bf k}\nu}\left|\beta^{I}_{m}\right\rangle.$ (22)
where $\lambda^{I}_{n{\bf k}\nu}$ are those calculated in Eq. 18.
2.3 Density functional perturbation theory
The phonon frequencies and displacement modes are obtained by diagonalization of the dynamical matrix:
$\omega_{\mathbf{q}}^{2}\mathbf{u}_{s\alpha}(\mathbf{q})=\sum_{s^{\prime}\beta}D_{s\alpha s^{\prime}\beta}(\mathbf{q})\mathbf{u}_{s^{\prime}\beta}(\mathbf{q}),$ (23)
where $D_{s\alpha s^{\prime}\beta}(\mathbf{q})$ is the dynamical matrix:
$D_{s\alpha s^{\prime}\beta}(\mathbf{q})=\frac{1}{\sqrt{M_{s}M_{s^{\prime}}}}\sum_{\nu}\frac{\partial^{2}E_{tot}}{\partial\mathbf{u}_{\mu s\alpha}\partial\mathbf{u}_{\nu s^{\prime}\beta}}e^{i\mathbf{q}\left(\mathbf{R}_{\nu}-\mathbf{R}_{\mu}\right)},$ (24)
where $E_{tot}$ is the DFT total energy, ${\bf q}$ is a wave vector in the Brillouin zone (BZ), $\mathbf{R}_{\mu}$ are the Bravais lattice vectors, $M_{s}$ are the atomic masses, and $\mathbf{u}_{\mu s\alpha}$ are the atomic displacements.
The second derivative of the DFT total energy can be written in terms of the change of the wave-functions due to a phonon perturbation projected on the conduction band. These functions are the solutions of a linear system: [45, 46]
$\left[H_{KS}^{{\bf k}+{\bf q}}+\alpha Q^{{\bf k}+{\bf q}}-\varepsilon_{{\bf k}\nu}S\right]P_{c}^{{\bf k}+{\bf q}}\frac{\partial u_{{\bf k}\nu}(\mathbf{r})}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}=-P_{c}^{{\bf k}+{\bf q}}\left[\frac{\partial V_{KS}}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}-\varepsilon_{{\bf k},\nu}\frac{\partial S}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}\right]u_{{\bf k}\nu}(\mathbf{r}),$ (25)
where $P_{c}^{{\bf k}+{\bf q}}$ is the projector in the conduction band and $\frac{\partial V_{KS}}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}=\frac{\partial V_{\text{loc }}}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}+\frac{\partial V_{H}}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}+\frac{\partial V_{\text{xc }}}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}+\frac{\partial V_{\text{NL }}}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}$. The change of the Hartree and exchange and correlation potential are:
$\frac{\partial V_{H}}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}$ $=\int\frac{e^{i{\bf q}({\bf r}^{\prime}-{\bf r})}}{\left|\mathbf{r}-\mathbf{r}^{\prime}\right|}\frac{\partial\rho\left(\mathbf{r}^{\prime}\right)}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}d^{3}r^{\prime},$ (26)
$\frac{\partial V_{xc}}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}$ $=\frac{dV_{xc}}{d\rho}\frac{\partial\rho(\mathbf{r})}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})},$
and depend self-consistently on the charge density induced by the perturbation:
$\frac{\partial\rho(\mathbf{r})}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}=4\sum_{{\bf k}\nu}\left[u_{{\bf k}\nu}^{*}(\mathbf{r})P_{c}^{{\bf k}+{\bf q}}\frac{\partial u_{{\bf k}\nu}(\mathbf{r})}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}\right]+\Delta_{{\bf u}_{s^{\prime}\beta}({\bf q})}({\bf r}),$ (27)
where the last term represents the change of the augmentation charge calculated in the ultrasoft and PAW case but not accelerated in the present work. [46] $Q^{{\bf k}+{\bf q}}$ is a projection in the valence manifold, [45] while the change of the nonlocal pseudopotential is described in more detail in the given references (see for instance Ref. [46]).
$\alpha Q^{{\bf k}+{\bf q}}$ can be written in the form:
$\alpha Q^{{\bf k}+{\bf q}}=\alpha\sum_{\mu}S|u_{{\bf k}+{\bf q}\mu}\rangle\langle u_{{\bf k}+{\bf q}\mu}|S,$ (28)
and its application to a set of wave functions $|x_{{\bf k}+{\bf q}\nu j}\rangle$ (here $j$ indicates the different perturbations ${s^{\prime}\beta}$) can be calculated easily using the fact that $S|x_{{\bf k}+{\bf q}\nu j}\rangle$ is already known from the routine that applies $H^{{\bf k}+{\bf q}}_{KS}$ and $S$. We have a first matrix-matrix multiplication:
$\mu_{{\bf k}+{\bf q}j\mu\nu}=\langle u_{{\bf k}+{\bf q}\mu}|S|x_{{\bf k}+{\bf q}\nu j}\rangle\alpha,$ (29)
then a second one:
$|y_{{\bf k}+{\bf q}\nu j}\rangle=\sum_{\mu}|u_{{\bf k}+{\bf q}\mu}\rangle\mu_{{\bf k}+{\bf q}j\mu\nu},$ (30)
and finally we must apply $S$ to the vectors $|y_{{\bf k}+{\bf q}\nu j}\rangle$ and we have:
$\alpha Q^{{\bf k}+{\bf q}}|x_{{\bf k}+{\bf q}\nu j}\rangle=S|y_{{\bf k}+{\bf q}\nu j}\rangle=|y_{{\bf k}+{\bf q}\nu j}\rangle+\sum_{Imn}q^{I}_{mn}|\beta^{I}_{m}\rangle\langle\beta^{I}_{n}|y_{{\bf k}+{\bf q}\nu j}\rangle,$ (31)
and this requires other three matrix-matrix multiplications as illustrated above.
The self-consistent linear system (Eq. 25) is solved by iterations. From an initial guess of the potentials, of ${\partial V_{H}\over{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}}+{\partial V_{xc}\over{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}}$, the linear system is solved and new induced charge and potentials are obtained. Mixing the latter with the potentials used in the linear system it is possible to reach a self-consistent solution.
The most time consuming step of this process is however the solution of the linear system at fixed ${\partial V_{H}\over{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}}+{\partial V_{xc}\over{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}}$ so we will focus on this step.
2.4 Preconditioned conjugate gradient
The algorithm used for the solution of Eq. 25 with a given right hand side is a preconditioned conjugate-gradient iterative algorithm. [47, 48, 49] Given a starting guess $x$ of the solution of the problem $Ax=b$, we improve it with the following algorithm:
$r$$\leftarrow$ $Ax-b,$ (32)
$d$$\leftarrow$ $M^{-1}r,$ (33)
$\rho$$\leftarrow$ $d^{T}r,$ (34)
$\gamma$$\leftarrow$ ${\rho\over\rho_{old}},$ (35)
$d$$\leftarrow$ $d+\gamma d_{old},$ (36)
$t$$\leftarrow$ $Ad,$ (37)
$\lambda$$\leftarrow$ $-{d^{T}r\over d^{T}t},$ (38)
$x$$\leftarrow$ $x+\lambda d,$ (39)
$r$$\leftarrow$ $r+\lambda t,$ (40)
$d_{old}$$\leftarrow$ $d,$ (41)
$\rho_{old}$$\leftarrow$ $\rho,$ (42)
and iterate from Eq. 33 until the modulus of $\rho$ is smaller than an input threshold. Here the arrows indicate that the variables on the left are substituted with those on the right. $r$ is the negative of the residual vector while $d$ contains minus the preconditioned residual in Eqs. 33 to 36 and minus the search direction from Eq. 36. Eqs. 35 and 36 are executed only from the second iteration onwards. The algorithm requires memory sufficient to save the vectors $r$, $d$, $t$, $d_{old}$ of the same size of the input vector $x$. $\rho$ and $\rho_{old}$, as well as $\gamma$ and $\lambda$, are instead scalars. Moreover, we need two external routines to apply $A$ and $M^{-1}$. The most time consuming step is the application of the matrix $A$ to $d$. In our case $A=H^{{\bf k}+{\bf q}}_{KS}+\alpha Q^{{\bf k}+{\bf q}}-\varepsilon_{{\bf k}\nu}S$ so again the acceleration rests on the routine that applies $H^{{\bf k}+{\bf q}}_{KS}$ and $S$ described above. For the preconditioning the following matrix diagonal in reciprocal space $M_{{\bf G},{\bf G}}=MAX(1.0,{|{\bf k}+{\bf q}+{\bf G}|^{2}\over 2\langle\psi_{{\bf k}+{\bf q}\nu}|-{1\over 2}\nabla^{2}|\psi_{{\bf k}+{\bf q}\nu}\rangle})$ is used and this vector is passed to the routine. The conjugate gradient algorithm is applied to each ${\bf k}$ point and to each $N_{pe}$ perturbations. The $N_{b}$ bands of a given ${\bf k}$ point are optimized together but the different k points and different perturbations are treated in sequence, one after the other. Only one ${\bf q}$ is calculated in each run. This algorithm has been used in the last thirty years in Quantum ESPRESSO to solve the linear system. Similar algorithms, with appropriate modifications, can be used also to minimize the total energy and solve the Kohn and Sham equations. [50]
3 GPU optimization
Figure 1: Algorithms used in the standard approach and in our optimized GPU approach for the diagonalization of the Hamiltonian.
Algorithm 1 CPU and standard GPU diagonalization
for ik = 1,nks do $\triangleright$ nks = #k points per pool
     build $H_{KS}^{\bf k}$ and $S^{\bf k}$
     compute $\varepsilon_{{\bf k}\nu}$ and $\psi_{{\bf k}\nu}$ by Davidson ($N_{b}$)
end for
Algorithm 2 Optimized GPU diagonalization
for ikb = 1,nkblock do $\triangleright$ nkblock = #k points blocks
     build in parallel $H_{KS}^{\bf k}$ and $S^{\bf k}$ on GPU threads ($N_{k}$)
     compute $\varepsilon_{{\bf k}\nu}$ and $\psi_{{\bf k}\nu}$ by Davidson on GPU threads ($N_{b}\times N_{k}$)
end for
Figure 2: Flowchart of the routine that applies $H^{\bf k}_{KS}$ and $S^{\bf k}$ to the wavefunctions. Close to each routine we write the number of threads that are used to run it on the GPU.
Quantum ESPRESSO has several levels of parallelization on the CPU. It is possible to divide the k-points in groups (called pools) and assign each group to a set of cores. These cores may be further divided in groups, with each group dealing with a set of bands (bands parallelization) and finally each group of cores dealing with a set of bands can further divide the reciprocal lattice vectors (G) and work only on a subset of these (G vectors parallelization). It is at this point that one can introduce the GPU acceleration. The standard method to use the GPU consists into allocating variables on the device memory (the GPU) and to call from HOST routines developed by NVidia that perform linear algebra (cuBLAS) operations or FFTs (cuFFT) on the data allocated on the device. It is also possible to add compiler directives to run loops in parallel on the GPU without changing the code.
The standard Davidson algorithm works sequentially on each ${\bf k}$ point of a pool and solves for all the bands (calling the routines that apply $H_{KS}$ and $S$ for a subset of bands, if bands parallelization is used). Therefore the number of times in which the GPU memory is loaded increases linearly with the number of ${\bf k}$ points. When the size of the problem is small, it can happen that the library matrix-matrix multiplications and FFTs routines cannot exploit all the capacity of the GPU because they have too few data to work on. As a result a GPU calculation might become even slower than a CPUs only calculation. Parallelizing on the ${\bf G}$ vectors just reduces further the size of the data allocated on the GPU for each k point and does not help in this case. Moreover, presently the GPU acceleration does not work well if many CPUs use the same GPU, so we use as many CPUs as GPUs.
Our strategy for accelerating the code on the GPU is illustrated schematically in Fig. 1. We put on the GPU memory as many wave-functions (i.e. k points) as possible in a block of $N_{k}$ k points and run simultaneously on all these k points the operations of the Davidson algorithm needed to diagonalize the Hamiltonian. Each pool of CPUs cores works on its set of k points as assigned by the pool parallelization of Quantum ESPRESSO and only these k points are divided in blocks for the GPU acceleration. Bands paralellization and G vector parallelization presently are not supported by our approach. The main GPU optimization has been performed on the routine that applies $H_{KS}$ and $S$ to the wave-functions $\psi_{{\bf k}\nu}$, but some acceleration has been obtained also carrying out the operations of the Davidson algorithm in parallel on many k points. In our approach, the routine that applies $H_{KS}$ and $S$ is a HOST routine (i.e. a routine running on the CPU) that receives as input $C_{{\bf k}+{\bf G}\nu}$ for $N_{k}$ k points, and gives as output the coefficients $C^{\prime}_{{\bf k}+{\bf G}\nu}$ and $C^{\prime\prime}_{{\bf k}+{\bf G}\nu}$ of the plane waves expansion of $H_{KS}\psi_{{\bf k}\nu}$ and of $S\psi_{{\bf k}\nu}$. This routine calls in sequence several GLOBAL routines (that is routines that run on the GPU and for which we can specify how many GPU threads run in parallel). The sequence of routines and the formula that they implement is illustrated in Fig. 2. The first computes the kinetic energy and runs $N_{pw}\times N_{k}\times N_{b}$ threads each one dealing with a ${\bf G}$ vector of one k point and of one band (here $N_{pw}$ is the number of ${\bf G}$ vectors used to expand the wave-functions). A second routine computes the scalar product in Eq. 18 and runs $N_{kb}\times N_{k}\times N_{b}$ threads, where $N_{kb}$ is the total number of projectors $|\beta^{I}_{m}\rangle$. The latter are loaded on the GPU for all the $N_{k}$ points before calling the Davidson algorithm. Another GLOBAL routine computes Eqs. 19 and 21 and runs $N_{k}\times N_{b}$ threads, while the sum over $n$ is made inside the routine. A routine copies $C_{{\bf k}+{\bf G}\nu}$ in $C^{\prime\prime}_{{\bf k}+{\bf G}\nu}$ and this is made in parallel running $N_{pw}\times N_{k}\times N_{b}$ threads. This is the first term of the application of $S$ to the wave functions. A routine sets to zero the FFT grid running $N_{\bf r}\times N_{k}\times N_{b}$ threads and another one sets the non zero elements of this grid running $N_{k}\times N_{b}$ threads, each one dealing with all the $N_{\bf r}$ grid points for one ${\bf k}$ point and one band. Then a set of three routines applies an inverse FFT to the wave-functions as detailed below, and a routine applies $V_{eff}$ running $N_{\bf r}\times N_{k}\times N_{b}$ threads. Another set of three routines applies the FFT to return to reciprocal space and a routine collects the results from the grid and adds them to $C^{\prime}_{{\bf k}+{\bf G}\nu}$. This is made in parallel on the GPU running $N_{k}\times N_{b}$ threads. Finally Eq. 20 is calculated by a routine that runs $N_{pw}\times N_{k}\times N_{b}$ threads and adds the result to $C^{\prime}_{{\bf k}+{\bf G}\nu}$. In the ultrasoft or PAW PPs case, the same routine calculates also the second term in the right hand side of Eq. 22 and adds it to $C^{\prime\prime}_{{\bf k}+{\bf G}\nu}$.
In this algorithm, $N_{k}$ must be carefully chosen and depends on the amount of GPU memory and on the size of the FFT grid. $N_{k}$ is mainly limited by the necessity to allocate on the GPU $N_{k}\times N_{b}$ FFT grids to apply, in parallel, the local potential to the Bloch functions. The allocation of this memory is done by the HOST routine that implements the Davidson algorithm.
We have also optimized some parts of the Davidson algorithm. The standard routine has been generalized introducing several loops on the $N_{k}$ k points and part of these loops have been transformed into kernel routines (GLOBAL routines) that perform the calculation in parallel using $N_{k}\times N_{b}$ threads. We have accelerated only the loops that took a significant amount of time. The other loops call the linear algebra cuBlas routines as in the standard approach.
Figure 3: Algorithms used in the standard phonon code and in our optimized GPU approach for solving the linear system that gives the perturbed wavefunctions.
Algorithm 3 CPU and standard GPU phonon algorithm
for ik = 1,nks do
     build $H_{KS}^{{\bf k}+{\bf q}}$ and $S^{{\bf k}+{\bf q}}$
     for ipert = 1,npe do $\triangleright$ npe = #perturbations
         build $P_{c}^{{\bf k}+{\bf q}}\frac{\partial V_{KS}}{\partial u_{s\alpha}({\bf q})}u_{{\bf k}v}$ ($N_{b}$)
         compute $\frac{\partial u_{{\bf k}v}}{\partial u_{s\alpha}({\bf q})}$ by conjugate gradient (CG) ($N_{b}$)
     end for
end for
Algorithm 4 Optimized GPU phonon algorithm
for ikb = 1,nkblock do $\triangleright$ $N_{k}$ = #k points per block
     build in parallel $H^{{\bf k}+{\bf q}}_{KS}$ and $S^{{\bf k}+{\bf q}}$ on GPU threads ($N_{k}$)
     build in parallel $P_{c}^{{\bf k}+{\bf q}}\frac{\partial V_{KS}}{\partial u_{s\alpha}({\bf q})}u_{{\bf k}v}$ on GPU threads ($N_{b}\times N_{k}\times N_{pe}$)
     compute $\frac{\partial u_{{\bf k}v}}{\partial u_{s\alpha}({\bf q})}$ by CG on GPU threads ($N_{b}\times N_{k}\times N_{pe}$)
end for
The acceleration of the phonon code instead has been carried out essentially on the algorithm that solves the linear system in Eq. 25. We proceed as in the Davidson algorithm (see the scheme in Fig. 3). However, in the phonon case the calculation of the induced charge density (Eq. 27) requires two FFT grids per band, one to contain $u_{{\bf k}\nu}^{*}(\mathbf{r})$ and one to contain $P_{c}^{{\bf k}+{\bf q}}\frac{\partial u_{{\bf k}\nu}(\mathbf{r})}{\partial{\bf u}_{s^{\prime}\beta}({\bf q})}$ so usually we use $N_{k}$ smaller than in the Davidson algorithm. The GPU optimization of the preconditioned conjugate gradient algorithm starts by allocating the COMPLEX vectors $g$, $d$, $d_{old}$, and $t$ on the GPU. For each variable $N_{k}\times N_{b}\times N_{pe}$ arrays are allocated. This memory is much larger than the one of the standard algorithm that requires only $N_{b}$ copies of each variable, but this space is allocated only on the GPU. The algorithm is then divided in loops over the $N_{k}$ k points, the $N_{pe}$ perturbations, and the $N_{b}$ bands. Loop one executes Eqs. 33 and Eqs. 34, loop two executes Eq. 36, and loop three computes $d^{T}r$ and $d^{T}t$ that appear in the numerator and denominator of Eq. 38. Finally loop four computes Eqs. 39, 40, and 41. Each loop is transformed into routine with the GLOBAL attribute that runs $N_{k}\times N_{b}\times N_{pe}$ threads, each one computing one perturbation to one band of one k point. Since each thread executes only a scalar product or an operation of the type $x\leftarrow x+\lambda d$ we have programmed these routines in CUDA Fortran without calling any other library routine. The array themselves instead are not split and each thread works on all the G vectors of each wave-function. All the other steps of the algorithm involve only scalar operations that are performed by the CPU.
Eq. 32 and Eq. 37 require an external routine to apply $A$ to the vectors $x$ (or $d$). For this operator we use the same routine that applies $H_{KS}$ and $S$ in the Davidson algorithm. The routine works in general for an arbitrary number of wavefunctions so when called from the conjugate gradient algorithm, in parallel on the GPU threads, it deals with the $N_{k}\times N_{pe}$ set of wavefunctions, each one composed by $N_{b}$ bands. We have then written a HOST routine that receives as input the coefficients $C^{\prime}_{{\bf k}+{\bf q}\nu j}$ and $C^{\prime\prime}_{{\bf k}+{\bf q}\nu j}$ of the Fourier transform of $H^{{\bf k}+{\bf q}}_{KS}|x_{{\bf k}+{\bf q}\nu j}\rangle$ and $S|x_{{\bf k}+{\bf q}\nu j}\rangle$ and gives as output the Fourier coefficients of $A|x_{{\bf k}+{\bf q}\nu j}\rangle$. This routine calls a series of GLOBAL routines for which we can control the number GPU threads that run in parallel. The first routine computes Eq. 29 and runs $N_{k}\times N_{pe}\times N_{b}\times N_{b}$ threads. A second routine computes $|a_{{\bf k}+{\bf q}\nu j}\rangle=H_{KS}^{{\bf k}+{\bf q}}|x_{{\bf k}+{\bf q}\nu j}\rangle-\varepsilon_{{\bf k}\nu}S|x_{{\bf k}+{\bf q}\nu j}\rangle$ and runs $N_{k}\times N_{pe}\times N_{b}$ threads. To complete the operator $A$ we have to calculate the operator $\alpha Q^{{\bf k}+{\bf q}}|x_{{\bf k}+{\bf q}\nu j}\rangle$ and we optimized also this part to run in many threads on the GPU in parallel on the ${\bf k}$ vectors, the bands, and the perturbations. This is done by calling another set of GLOBAL routines. The first computes Eq. 30 running on $N_{k}\times N_{pe}\times N_{b}$ threads, another one computes the scalar products $\langle\beta^{I}_{n}|y_{{\bf k}+{\bf q}\nu j}\rangle$ that appear in Eq. 31 and runs $N_{k}\times N_{pe}\times N_{kb}\times N_{b}$ threads, and a third routine calculates Eq. 21 using the scalar products just calculated and runs on $N_{k}\times N_{pe}\times N_{b}$ threads. Finally a GLOBAL routine computes Eq. 31 and adds it to $|a_{{\bf k}+{\bf q}\nu j}\rangle$ running in $N_{k}\times N_{pe}\times N_{b}\times N_{pw}$ threads.
4 Fast Fourier transform
The application of $V_{eff}({\bf r})$ to one Bloch wave-function requires two Fourier transforms. It is convenient to introduce a mesh in reciprocal space:
${\bf G}_{m_{1},m_{2},m_{3}}\equiv m_{1}{\bf b}_{1}+m_{2}{\bf b}_{2}+m_{3}{\bf b}_{3},$ (43)
where ${\bf b}_{1}$, ${\bf b}_{2}$, and ${\bf b}_{3}$ are the principal reciprocal lattice vectors and $m_{1}$, $m_{2}$, and $m_{3}$ are integers, and a mesh in real space:
${\bf r}_{l_{1},l_{2},l_{3}}={l_{1}\over N_{1}}{\bf a}_{1}+{l_{2}\over N_{2}}{\bf a}_{2}+{l_{3}\over N_{3}}{\bf a}_{3},$ (44)
where ${\bf a}_{1}$, ${\bf a}_{2}$, and ${\bf a}_{3}$ are the direct lattice vectors, and $l_{1}$, $l_{2}$, and $l_{3}$ are integers. The integers $N_{1}$, $N_{2}$, and $N_{3}$ define the size of the mesh in real space and, equivalently, the size of the mesh in reciprocal space. They must be sufficiently large so that the vectors ${\bf G}_{m_{1},m_{2},m_{3}}$ contain all the vectors ${\bf G}-{\bf G}^{\prime}$ defined by the basis set.
[…]

### v2

@@
 journal: Computer Physics Communications
+The published version of this paper can be found at DOI:
+https://doi.org/10.1016j.cpc.2024.109439
 1 Introduction
@@
 They must be sufficiently large so that the vectors ${\bf G}_{m_{1},m_{2},m_{3}}$ contain all the vectors ${\bf G}-{\bf G}^{\prime}$ defined by the basis set.
-[…]
+Given a function in reciprocal space, defined on the G vectors $\tilde{f}(m_{1},m_{2},m_{3})\equiv f({\bf G}_{m_{1},m_{2},m_{3}})$, its real space form $f(l_{1},l_{2},l_{3})=f({\bf r}_{l_{1},l_{2},l_{3}})$ is given by:
+$f(l_{1},l_{2},l_{3})=\sum_{m_{1}=0}^{N_{1}-1}\sum_{m_{2}=0}^{N_{2}-1}\sum_{m_{3}=0}^{N_{3}-1}\tilde{f}(m_{1},m_{2},m_{3})e^{i2\pi l_{1}m_{1}/N_{1}}e^{i2\pi l_{2}m_{2}/N_{2}}e^{i2\pi l_{3}m_{3}/N_{3}}.$ (45)
+The transform is made in three steps. In the first step, we compute $N_{1}\times N_{2}$ one dimensional FFTs along $z$:
+$\bar{f}(m_{1},m_{2},l_{3})=\sum_{m_{3}=0}^{N_{3}-1}\tilde{f}(m_{1},m_{2},m_{3})e^{i2\pi l_{3}m_{3}/N_{3}}.$ (46)
+We run $N_{1}\times N_{b}\times N_{k}$ threads on the GPU by calling a GLOBAL routine, and each thread computes $N_{2}$ FFTs.
+Each FFT (sum over $m_{3}$) is carried out by calling an FFT library routine (cfft1b from fftpack.5.1) which is declared as a DEVICE routine.
+In the second step, we compute:
+$\hat{f}(m_{1},l_{2},l_{3})=\sum_{m_{2}=0}^{N_{2}-1}\bar{f}(m_{1},m_{2},l_{3})e^{i2\pi l_{2}m_{2}/N_{2}}.$ (47)
+In this case we run $N_{1}\times N_{b}\times N_{k}$ threads each one doing $N_{3}$ FFTs.
+Each FFT (sum over $m_{2}$) is carried out by the DEVICE FFT library routine cfft1b.
+Finally, to complete the three dimensional Fourier transform, in the third step we calculate:
+$f(l_{1},l_{2},l_{3})=\sum_{m_{1}=0}^{N_{1}-1}\hat{f}(m_{1},l_{2},l_{3})e^{i2\pi l_{1}m_{1}/N_{1}}.$ (48)
+In this case we run $N_{k}\times N_{b}$ threads on the GPU each one computing $N_{2}\times N_{3}$ FFTs.
+Each one dimensional FFT (sum over $m_{1}$) is carried out by the DEVICE FFT library routine cfft1b.
+In a similar way one can make a three dimensional Fourier transform to obtain the reciprocal space function from its real space form:
+$\tilde{f}(m_{1},m_{2},m_{3})={1\over N_{\bf r}}\sum_{l_{1}=0}^{N_{1}-1}\sum_{l_{2}=0}^{N_{2}-1}\sum_{l_{3}=0}^{N_{3}-1}f(l_{1},l_{2},l_{3})e^{-i2\pi l_{1}m_{1}/N_{1}}e^{-i2\pi l_{2}m_{2}/N_{2}}e^{-i2\pi l_{3}m_{3}/N_{3}},$ (49)
+where $N_{\bf r}=N_{1}N_{2}N_{3}$. In this case we call the DEVICE function cfft1f to actually carry out the one dimensional FFTs.
+The product of $V_{eff}$ with the wave-function:
+$u_{{\bf k}\nu}^{\prime}(l_{1},l_{2},l_{3})=V_{eff}(l_{1},l_{2},l_{3})u_{{\bf k}\nu}(l_{1},l_{2},l_{3}),$ (50)
+is made by running $N_{\bf r}\times N_{k}\times N_{b}$ threads on the GPU, each thread computing one product.
+After computing the product, an FFT as in Eq.
+49 gives the Fourier components of the product that can be added to those obtained by applying the kinetic energy.
+This FFT is performed by three routines similar to those described for the inverse FFT.
+4.1 FFT on the device
+Eqs.
+46,47,48 cannot be implemented as written since they involve $N_{i}^{2}$ operations, where $N_{i}$ is $N_{1}$, $N_{2}$ or $N_{3}$.
+These sums can be done more efficiently with an FFT algorithm that requires $N_{i}log(N_{i})$ operations.
+[47] The FFTXlib of Quantum ESPRESSO contains both the three dimensional FFT driver and a copy of an old FFTW library.
+[51] It also supports the newer FFTW3 library, some vendor-specific FFT libraries, and it can call library routines optimized for the GPU in cuFFT [9, 20].
+Moreover, it can carry out the FFT in parallel when the FFT mesh (and G vectors) are distributed among different MPI processes.
+However, these routines are called from the CPU HOST with actual argument variables that are allocated on the GPU and they take care of launching the kernel threads on the GPU.
+In our approach, the FFT routines are called from inside the GPU threads and therefore must have the DEVICE attribute, hence FFTXlib cannot be used.
+The library that offers this functionality cuFFTDx is written in C++ and it has not yet a FORTRAN interface.
+Therefore, we have taken the fftpack5.1 [52] which is distributed under the GNU GPL licence together with its Fortran source and we have modified each routine and function of this library by adding the ATTRIBUTES(DEVICE).
+We have also constructed a Fortran interface for each routine so that the routines that include the interface can know that the routines of fftpack are actually DEVICE routines and accept variables allocated on the GPU.
+The modified library is distributed together with the thermo_pw package.
+5 Matrix Diagonalization
+The Davidson algorithm requires the solution of a generalized eigenvalue problem in a reduced basis:
+$Ax=\lambda Bx,$ (51)
+where $A$ and $B$ are Hermitian matrices and $\lambda$ and $x$ are the eigenvalues and eigenvectors.
+Usually, the CPU makes this calculation by calling LAPACK routines [43] such as ZHEGVX that computes selected eigenvalues and, optionally, eigenvectors of a complex generalized Hermitian-definite eigenproblem, or ZHEGV that computes all eigenvalues and eigenvectors of the same matrices.
+It is also possible to call library routines optimized for the GPU of the cuSolver or of the MAGMA libraries.
+A HOST driver that calls these routines is contained in the LAXlib library distributed with Quantum ESPRESSO.
+We have tested this approach creating a loop over the $N_{k}$ k points that calls these routines, but found that it is possible to obtain a significant speed up by simultaneously diagonalizing the generalized eigenvalue problem for many k points.
+We run therefore a GLOBAL routine with as many threads as possible (ideally $N_{k}$, but see below).
+In order to solve the generalized eigenvalue problem inside a GLOBAL routine we cannot call HOST routines such as those available in cuSolver or in MAGMA [21] what is needed is a library that can be called from inside the GPU threads (with DEVICE routines).
+Since we are not aware of any DEVICE implementation of LAPACK, we took the routines ZHEGVX and ZHEGV together with those called by them, transformed them into DEVICE routines, and wrote the corresponding Fortran interfaces.
+We found only one problem with this approach: The routine ZPOTRF2, which performs the Cholesky factorization of a Hermitian positive definite matrix $A$, is recursive.
+Since CUDA Fortran does not allow for recursive DEVICE routines or functions, we rewrote it with a non recursive algorithm.
+The number of k points that can be diagonalized simultaneously is usually lower than $N_{k}$ since the LAPACK DEVICE routines use a certain amount of GPU resources.
+So we divided the $N_{k}$ k points in blocks of maximum size determined empirically on the available machine.
+6 Results
+6.1 Benchmark Example
+We have implemented our approach in the thermo_pw code [53] which is a driver of Quantum ESPRESSO routines to calculate materials properties.
+To activate the new approach, it suffices to set the flag many_k to .TRUE.
+and the input variable memgpu to the amount of GPU memory (in GBytes).
+Both variables are written in the thermo_control input file.
+The new routines are in the directory qe of thermo_pw, while the LAPACK and fftpack5.1 routines modified with the ATTRIBUTES(DEVICE) together with their interfaces are distributed in separate subdirectories of the thermo_pw package.
+For further details please refer to the thermo_pw user’s guide.
+Our benchmark is a part of the calculations carried out to compute the quasi-harmonic temperature dependent elastic constants of tungsten.
+[29] Our system is body centered cubic (bcc) tungsten simulated with the PBEsol exchange and correlation functional [54] at the lattice constant $a=5.965$ a.u..
+Tungsten is described with a PAW pseudopotential that has 141414 valence electrons and we compute $N_{b}=11$ bands.
+[55] We use cut-offs for the wave-functions/charge density of 909090/360360360 Ry, a k-point mesh of $\times 45\times 45$ 45 and deal with the Fermi surface with the smearing approach ([28]) with a smearing parameter $\sigma=0.02$ Ry.
+The FFT mesh has size $\times 32\times 32$ 32 for a total of $N_{\bf r}=32768$ mesh points.
+We compute the phonon frequencies for the point ${\bf q}={2\pi\over a}(-1/8,-1/4,3/8)$.
+The small space group of this ${\bf q}$ point has no rotational symmetry in it, so we need to use the complete mesh of $45^{3}=91125$ ${\bf k}$ points when computing the perturbed wave-functions.
+Since we need also the eigenvalues and eigenfunctions at ${\bf k}+{\bf q}$ we compute the band structure of 182250182250182250 ${\bf k}$ points.
+We report the time obtained with version 7.3 of Quantum ESPRESSO together with thermo_pw version 2.0.0.
+All tests have been performed on the Leonardo supercomputer at CINECA.
+Each node of the machine has a CPU with 323232 cores and 444 Ampère GPUs.
+In the Leonardo manual, the theoretically declared peak performance of one node (323232 cores) is 168016801680 Gflops while the four GPUs of one node can provide 750007500075000 Gflops.
+There is therefore a maximum theoretical acceleration of a factor of 454545.
+We run on the GPUs using as many CPU cores as GPUs and each CPU runs one MPI process.
+[56] MPI processes can communicate among themselves with MPI library calls.
+Each MPI process communicates with one GPU, multiple MPI processes using the same GPU are not allowed.
+Moreover we do not use direct GPU-GPU communication.
+When several MPI processes run, the total number of k points is divided in a number of pools equal to the number of MPI processes.
+The code is compiled with the PGI Fortran compiler contained in the Nvidia SDK [57].
+6.2 FFT
+Table 1: Comparison of the time spent by computing the FFT and the inverse FFT when applying the Hamiltonian operator in the Davidson algorithm and in the conjugate gradient algorithm for the example described in the paper.
+CPU GPU optimized GPU
+#CPU 32 32 32 64 64 1 2 4 8 1 2 4 8
+#GPU 0 0 0 0 0 1 2 4 8 1 2 4 8
+#task(np) 32 32 32 64 64 1 2 4 8 1 2 4 8
+#pool(nk) 8 16 32 32 64 1 2 4 8 1 2 4 8
+time (s) 6771 6280 4539 3269 2213 13415 6866 3441 1744 7532 3574 2068 934
+In Table 1, we report the time necessary to compute the FFTs to apply the local potential.
+We consider three cases: CPUs only, standard GPU code that calls the cuFFT library, and the optimized GPU code that uses the fftpack.5.1 routines declared as DEVICE routines.
+In the GPUs runs, we consider 111, 222, 444 or 888 GPUs.
+For the CPUs only runs, we use all the CPUs of one (323232) or two nodes (646464).
+Further, with 323232 cores, the k points are divided into 888, 161616, or 323232 pools, with 646464 cores, into 323232 or 646464 pools.
+When comparing CPUs and GPUs, we compare 444 or 888 GPUs with the best times obtained with 323232 or 646464 cores, respectively.
+We start by discussing the CPUs only case.
+With both 323232 or 646464 cores, the minimum FFT time is obtained when the number of pools is equal to the number of cores.
+This indicates that in this system it is not useful to divide the ${\bf G}$ vectors among CPUs.
+The second observation is that when we pass from one to two nodes the time halves, showing a good scaling with the number of nodes.
+We call $T_{cpu}$ the best time obtained with one or two nodes.
+Passing now to the GPU times, we see that both with the standard algorithm and with the optimized one the computational time is inversely proportional to the number of GPUs.
+Comparing now the time taken by the standard GPU algorithm, we see that it is 0.760.760.76 Tcpu (444 GPUs), 0.790.790.79 Tcpu (888 GPUs).
+So, as far as the FFT is concerned, it is convenient to use the GPUs instead of the CPUs although the gain is not big.
+The optimized GPU algorithm gives times that are 0.460.460.46 Tcpu (444 GPU), 0.420.420.42 Tcpu (888 GPUs).
+This is much less than the theoretical capacity of the GPU, but still it makes convenient to use the latter.
+When computing the FFT, the optimized GPU algorithm is 1.91.91.9 times faster (888 GPUs) than the standard GPU version that calls the CuFFT routines in sequence on the k points.
+6.3 Diagonalization
+Table 2: Comparison of the time spent to diagonalize the reduced Hamiltonian using linear algebra routines (within the Davidson algorithm).
+CPU GPU optimized GPU
+#CPU 32 32 32 64 64 1 2 4 8 1 2 4 8
+#GPU 0 0 0 0 0 1 2 4 8 1 2 4 8
+#task(np) 32 32 32 64 64 1 2 4 8 1 2 4 8
+#pool(nk) 8 16 32 32 64 1 2 4 8 1 2 4 8
+time (s) 81 44 20 21 10 1286 684 348 178 220 118 60 29
+In Table 2, we report the time spent by the diagonalization of the reduced Hamiltonian carried out by the LAPACK routines on the CPU, by the cuSolver library running on GPU called by the LAXlib package, and by the optimized GPU version of the code in which the Hamiltonians of many k-vectors are diagonalized simultaneously by the LAPACK routines declared as DEVICE routines.
+The effect of using pools and several CPUs is also illustrated.
+The sizes of the matrices to be diagonalized vary, depending on the istantaneous size of the basis set in the Davidson algorithm.
+The routine must find the lowests $N_{b}=11$ eigenpairs in a matrix that can have a maximum size equal to $4N_{b}=44$.
+This is repeated for all k points for all Davidson iterations and all self-consistent iterations in addition to a band structure calculation before the phonon calculation (in which there are about $2\times 10^{5}$ k points).
+The CPU diagonalization time scales linearly with the number of ${\bf k}$ points and therefore depends only on the number of pools.
+Using 646464 or 323232 cores gives exactly the same time when we use 323232 pools, but if we use a number of pools equal to the number of cores with two nodes we halves the diagonalization time with respect to one node.
+A good scaling is also shown by the GPU calculation.
+Increasing the number of GPUs increases the number of pools and therefore decreases the number of ${\bf k}$ points per pool.
+With both the standard GPU algorithm and with the optimized one we could not run faster than the CPU.
+With the standard algorithm the size of the matrix to diagonalize is so small that the time to initialize the GPU greatly exceeds the CPU diagonalization time.
+In this particular example, the time of the standard GPU calculation is 181818 Tcpu.
+With our optimization we could reduce this time to 333 Tcpu.
+In our example however the total time for the diagonalization is small with respect to all other times and we have not tried to further optimize this part.
+6.4 Application of the Hamiltonian and of S
+Table 3: Comparison of the total time spent to apply $H_{KS}$ and $S$ to the wave-functions in the Davidson algorithm and in the conjugate gradient algorithm.
+CPU GPU optimized GPU
+#CPU 32 32 32 64 64 1 2 4 8 1 2 4 8
+#GPU 0 0 0 0 0 1 2 4 8 1 2 4 8
+#task(np) 32 32 32 64 64 1 2 4 8 1 2 4 8
+#pool(nk) 8 16 32 32 64 1 2 4 8 1 2 4 8
+time (s) 7907 7319 5519 3780 2702 36157 18465 9263 4694 9138 4375 2472 1138
+time -timeFFT (s) 1136 1039 980 511 489 22742 11599 5822 2950 1606 801 404 204
+In Table 3 we show the time required for the application of the Hamiltonian and of the overlap matrix $S$ to the wave-functions.
+This time comprises the time needed to apply the FFT and inverse FFT to the wave-functions, the time needed to apply the kinetic energy and the nonlocal pseudopotential as well as the time needed to apply the overlap matrix $S$.
+In the same table we report also the difference between these times and the times needed to carry out the FFT reported in Table 1.
+In the time reported in the table, we apply the operator $H_{KS}$ and $S$ about $8\times 10^{7}$ times (as reported by the code when we do not use the optimized algorithm and $7.3\times 10^{4}$ when we use the optimized algorithm and many k-points are calculated concurrently).
+This is reasonable since we have $1\times 10^{5}$ k points, about 161616 self-consistent iterations and $N_{pe}=3$ modes.
+This gives an average of 161616 conjugate gradient steps per iteration.
+To count the number of operations is more difficult since the number of bands is not always constant.
+If take as an average value $N_{b}=11$ bands, the number of plane waves $N_{pw}=2093$ and a number of projector functions $N_{kb}=18$ we see that Eq.
+38 is the multiplication of a matrix 18×202318$\times$ 202318 × 2023 and a matrix 2023×112023$\times$ 112023 × 11.
+We start by considering the CPU times when FFT time is subtracted.
+These times depend on the number of cores, but less on how these cores are distributed between ${\bf G}$ vectors and ${\bf k}$-point pools.
+Still using only ${\bf k}$-point pools gives the shortest times but the differences are small.
+Comparing with the standard GPU version, we see that the application of the nonlocal potential and of the $S$ matrix require too many small size matrix-matrix multiplications and this part of the calculation is quite slow on the GPU.
+For this calculation the required time is 666 Tcpu.
+The optimized GPU algorithm is much faster and needs about 0.420.420.42 Tcpu.
+Adding also the speedup obtained with the FFT, the optimized GPU algorithm takes about 0.440.440.44 Tcpu.
+Comparing the two GPU algorithms, the optimized one is 444 times faster in applying $H_{KS}$ and 141414 times faster in applying the nonlocal pseudopotential and the $S$ matrix.
+6.5 Total time
+Table 4: Total time spent in the standard CPU calculations.
+The number of CPUs, GPUs, tasks, and pools are also indicated.
+The number of core-hours is obtained multiplying the total time by the number of cores.
+CPU
+#CPU 32 32 32 64 64 128 256
+#GPU 0 0 0 0 0 0 0
+#task(np) 32 32 32 64 64 128 256
+#pool(nk) 8 16 32 32 64 128 256
+time (s) 11400 10560 8040 5640 4500 2280 1182
+time (m) 190 176 134 94 75 38 20
+core-hours 101 94 71 100 80 81 85
+Table 5: Total time spent in the standard GPU calculations.
+This includes also the time passed on the part of the code that are not GPU accelerated or are not GPU optimized.
+The number of CPUs, GPUs, tasks, and pools are equal.
+GPU
+#CPU 1 2 4 8 16 32
+time (s) 107640 54780 27720 14100 7500 3720
+time (m) 1794 913 462 235 125 62
+core-hours 239 243 246 250 267 265
+Table 6: Total time spent in the optimized GPU calculations.
+This includes also the time passed on the part of the code that are not GPU accelerated or are not GPU optimized.
+The number of CPUs, GPUs, tasks, and pools are equal.
+optimized GPU
+#CPU 1 2 4 8 16 32
+time (s) 16080 7860 4680 2153 1140 585
+time (m) 268 131 78 36 19 10
+core-hours 36 35 42 38 41 43
+In this section we present some benchmarks of the entire run, considering both the self consistent and the phonon frequencies calculations.
+We report in Tables 456 the total time.
+This time is approximately twice the time required by the application of $H_{KS}$ and $S$ in the CPU and in the optimized GPU cases and three times in the standard GPU case.
+Considering now the total Tcpu, we see that the faster time is obtained when the number of pools is equal to the number of cores.
+The scaling with the number of nodes is good: from 323232 to 646464 cores the code is 1.81.81.8 times faster.
+The standard GPU approach takes 3.43.43.4 Tcpu (one node) or 3.13.13.1 Tcpu (two nodes), while the optimized GPU approach takes 0.580.580.58 Tcpu (one node) and 0.480.480.48 Tcpu (two nodes).
+The difference between one and two nodes is due to the different number of CPUs cores available in the two cases.
+The parts that are not accelerated are calculated faster when more cores are available.
+Comparing now the two GPU algorithms we see that the optimized one is about 666 times faster.
+In the table we have indicated also the cost of each run in core-hours.
+This cost is obtained by multiplying the total time by the number of core used (in the GPU case, each GPU costs 888 cores).
+We have also added the time needed with 444 and 888 nodes (128128128 and 256256256 cores).
+We find that increasing the number of nodes the total cost tend to increase (even if there are some fluctuations) since it is difficult to achieve an exact linear scaling with the number of pools.
+In the optimized GPU case the optimum is obtained with 222 nodes.
+It is therefore convenient to carry out this calculations with a small number of nodes per ${\bf q}$ point and calculate in parallel on different nodes different ${\bf q}$ points and geometries.
+However, even with an ideal scaling with the pools and a computer that can provide as many GPUs as desired, it is still convenient to use pools that contain a number of ${\bf k}$ point sufficient to occupy the GPU memory and use it (gaining about a factor 2X), than split the calculations so that each pool has a single k point.
+7 Conclusions and perspectives
+We discussed a scheme to accelerate on the GPUs electronic structure codes based on plane waves and pseudopotentials.
+We have shown in the example of bcc tungsten that our scheme can be faster than the currently implemented GPU version when the system has small unit cells but requires a thick mesh of k points.
+The main idea is to apply the Hamiltonian to the wave-functions in parallel on many ${\bf k}$ points, one per GPU thread, so as to increase both the size of the data on which the GPU works at any given time and to give to the GPU a sufficient numerical workload to exploit all its SMs.
+Our method has been implemented in CUDA Fortran by partially rewriting the code and by using GLOBAL and DEVICE routines to parallelize the work of different GPU threads.
+We have discussed in detail the optimization of the Davidson algorithm, the application of the Kohn and Sham Hamiltonian and of the overlap matrix $S$ to the wave-functions, and the preconditioned conjugate gradient algorithm which is used to solve the linear system of DFPT.
+In our example the application of $H_{KS}$ and $S$ to the wave-functions accounts for about one half of the total time with CPUs and about 1/31/31 / 3 with the GPUs.
+For these operations our optimized GPU method is about 666 times faster than the standard GPU approach, and about twice as fast than the CPUs only calculation.
+The main limitation of the present implementation is that it does not support the reciprocal lattice vectors distribution among CPUs.
+It is instead possible to divide the k points in pools so that different GPUs acts on different pools.
+Finally, we underline the fact that when the system (and the FFT mesh) becomes large enough the cuFFT library routines become more efficient than our DEVICE routines and at that point the standard approach might become more convenient.
+Our approach required a precise control of the GPU threads and math libraries (with DEVICE functions) that can be called from inside the GPU threads.
+Presently not many libraries offer this functionality and we hope that, in future, optimized DEVICE versions of math libraries will appear together with FORTRAN compatible interfaces.
+The substitution of our transformed routines with better optimized ones could further improve the speed of our code.
+As a last consideration we might ask if there are other ways to speed up the plane-waves pseudopotential codes for metallic cases as those that we need for our research.
+There are several option that one might explore from introducing a batched form of the FFT and of the linear algebra routines, to using new FFT GPU libraries such as heFFTe[58].
+If these options would solve the problem pointed out in this paper within the standard GPU scheme remains still to be investigated.
+The implemented software is distributed within the GPL licence within the thermo_pw package. [53]
+8 Appendix: GPU and CUDA Fortran
+The CUDA architecture is built around a scalable array of multithreaded Streaming Multiprocessors (SMs).
+Each SM has a set of execution units and a set of registers and can operate on variables contained in the GPU memory.
+CUDA Fortran allows the allocation of data on the GPU (called DEVICE in this context), the transfer of data from and to the GPU and the writing of routines that can run on the GPU in many different threads, each one working on different data (called GLOBAL) or that can be called from the GPU threads (called DEVICE).
+This is possible also with OpenACC and openMP compiler directives but we have opted for CUDA Fortran since in this moment there is a large basis of installed supercomputers equipped with NVIDIA GPUs that can run code written in CUDA Fortran and also the one available to us is in this category.
+The use of OpenACC and openMP that could be required to make our code transferable to GPUs of other vendors might be considered in the future if necessary.
+In CUDA Fortran, to run on the GPU, one declares the routines with ATTRIBUTES(GLOBAL) or ATTRIBUTES(DEVICE).
+Both run on the GPU, but the first can be called from the CPU host with the triple chevron syntax ($<$,$>$)($<$,$>$)( < < < , > > > ) to specify the number of threads blocks and threads per block that are employed.
+In general thread blocks can be arranged in a three dimensional grid with variable size in each dimension.
+The second can be called from the GLOBAL routines on the data already selected for the current thread.
+CUDA makes four pieces of information available to each thread: the thread index (threadIdx), the block index (blockIdx), the size and shape of a block (blockDim), and the size and shape of a grid (gridDim).
+This information can be used to choose the variables the current thread will work on.
+To give an order of magnitude, the Volta (Ampère) GPU architecture has 848484 (108108108) SMs each capable of running up to 323232 threads that is 32×84=268832$\times$ 84=268832 × 84 = 2688 (345634563456) threads can run simultaneously on the GPU, however each thread block must run the same instructions on different data, while different thread blocks can execute different instructions.
+The code is independent from the GPU architecture on which it will run and can require even bigger grids and block sizes whose threads are run in sequence on the available SMs.
+9 Acknowledgments
+This work has been supported by the Italian MUR (Ministry of University and Research) through the National Centre for HPC, Big Data, and Quantum Computing (grant No.
+CN00000013).
+Computational facilities have been provided by SISSA through its Linux Cluster, ITCS, and the SISSA-CINECA 2021-2024 Agreement.
+Partial support has been received by MAX “MAterials design at the eXascale” Centre of Excellence for Supercomputing applications (Grant agreement No.
+101093374, co-funded by the European High Performance Computing joint Undertaking (JU) of the European Union and participating countries).
+X.
+Gong acknowledges the support received in the framework of the Joint Research Agreement for Magnetic Confinement Fusion between Eni and CNR.